3 years ago, larry was twice as old as ben. three years from now hewill be 1 1/2 times as old as ben. how old is each?

Answer by ewatrrr(24785)


Hi,
Let x be ben's age now and y represent larry's age now.
y - 3 =2(x -3)=2x - 6 3 years ago, larry was twice as old as ben
y + 3 =(3/2)(x + 3)=(3/2)x + (9/2) three years from now he will be 1 1/2 times as old as ben
y - 3 =2x - 6
y + 3 =(3/2)x + (9/2)
Solving for x by eliminating y
y - 3 =2x - 6
-y - 3 =-(3/2)x - (9/2) subtracting second equation from the first
-6 = x/2 - 21/2
9/2 = x/2
9 = x, ben's age now.
Substituting for x to solve for y, larry's age now
y - 3 = 2x - 6
y = 2x -3
y =2*9 - 3 = 15, larry's age now
checking our answer
3 years ago, larry was twice as old as ben
12 = 2*6